#3 Local groundwater has been contaminated with three industrial solvents. We wish to treat 1500 gpm of the water using air stripping in a column with 10 stages. What is vour recommendation for the required air flow rate (in SCFM) to produce water that meets EPA standards. Required data are given below Compound K-Value |Concentration in EPA Maximum Allowed Groundwater (ppm) Concentration (ppm 1.2-Dichloroethane Trichloroethylene 「·「·loroethane Also calculate the concentration of each of the organic species in the cleaned water. The concentrations are in ppm by weight. 60 650 275 85 120 145 0.005 0.005 0.200 1,1,1-Trichloroethane|
Answer #1

For any component, Air flowrate = K*(PPM,in-PPM,out)*10^-6*flowrate(gpm)*.13 SCFM

=>For 1,2 Dichloroethane, Airflowrate = 60*(85-.005)*10^-6*1500*.13 =0.99 SCFM

=>For Trichloroethylene, Air flowrate = 650*(120-.005)*10^-6*1500*.13 =15.2 SCFM

=> For 1,1,1 Trichloroethane, Air flowrate = 275*(145-.2)*10^-6*1500*.13 =7.76 SCFM

Therefore, Air flowrate required is maximum of all three = 15.2 SCFM

=> Concentrations of 1,2 Dichloroethane in Cleaned water = 0 ppm

=>Concentrations of Trichloroethylene in Cleaned water = .005 ppm

=>Concentrations of 1,1,1 Trichloroethane in Cleaned water = 0 ppm